OpenStudy (anonymous):

21 is 30% of what number?

OpenStudy (anonymous):

.30*N = 21

OpenStudy (anonymous):

30 / 100 * x = 21 30 x = 2100 x = 70

OpenStudy (anonymous):

THANK YOU BT HOW U DID IT

OpenStudy (anonymous):

it's clear enough.....30% * x = 21 then multiply 21 with 100 2100 then divide it to 30 and the answer is 70

OpenStudy (anonymous):

70

OpenStudy (anonymous):

saad saad is right,...

OpenStudy (anonymous):

alternative/longwinded version of tian .30*n=21 divide both sides by .30 .30/.30 give you 1 so the left is 1N or just N 21/.30 = 70 so N = 70

OpenStudy (anonymous):

yeah U're rught jtho